OpenStudy (anonymous):

Evaluate the integral of: du/5√u

OpenStudy (anonymous):

\[\frac{du }{ 5\sqrt{u}}\]

geerky42 (geerky42):

Hint: \[\large\int\limits x^n dx = \frac{ x^{n+1} }{ n+1 }\]

zepdrix (zepdrix):

\[\Large\rm \int\limits\limits \frac{du}{5 \sqrt u}=\int\limits\limits\frac{1}{5\sqrt u}du=\frac{1}{5}\int\limits\limits \frac{1}{\sqrt u}du=\frac{1}{5}\int\limits u^{-1/2}du\]From that point, apply the power rule that geerky described.